OpenStudy (anonymous):

adam has 4 dimes, 4 quarters and 4 nickels in his pocket. He pulls out 2 coins making sure to put the first one back in his pocket before picking the second. What are the odds that he pulls out 2 nickels? 1/3 1/9 1/16 1/81

terenzreignz (terenzreignz):

The odds of picking a nickel in the first pull is 1/3, as there are 4 nickels and 12 coins. the odds of picking a nickel in the second pull is 1/3, as there will be the same number of nickels against coins (4 to 12) The odds of picking two nickels is the product of 1/3 and 1/3 which is...? :)

OpenStudy (farmdawgnation):

Sorry guys. I'll fix that.

OpenStudy (anonymous):

probability he gets two nickels with replacement is \(\frac{1}{3}\times \frac{1}{3}=\frac{1}{9}\) but ODDS are one to eight

OpenStudy (anonymous):

none of your possible answers appear correct

OpenStudy (anonymous):

careful here odds not the same number as probability